OpenStudy (anonymous):

Two objects, each with a charge of +5.0 × 10^-5 C, are separated by 0.50 m. What is the magnitude and type of force between them? A. -90 N, attractive B. -45 N, repulsive C. +45 N, attractive D. +90 N, repulsive

OpenStudy (michele_laino):

please wait a moment

OpenStudy (anonymous):

ok!

OpenStudy (michele_laino):

the force is repulsive, since both charges are positive

OpenStudy (anonymous):

ok! so it is either B or D? how can we find out which one?

OpenStudy (michele_laino):

the magnitude of that force, is given by the Coulomb's law: \[\large F = K\frac{{{Q_1}{Q_2}}}{{{d^2}}} = 9 \times {10^9}\frac{{5 \times {{10}^{ - 5}} \times 5 \times {{10}^{ - 5}}}}{{{{\left( {5 \times {{10}^{ - 1}}} \right)}^2}}}\]

OpenStudy (anonymous):

ok! we get 90? so our solution is choice D?

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

yay!! thank you!!

OpenStudy (michele_laino):

:)